OpenStudy (anonymous):

The equation tan(x-pi/3) is equal to _____?

OpenStudy (anonymous):

@zepdrix can you help me please?

OpenStudy (mathmale):

Hello, Jules, Remember that the tangent function is defined in this following manner (among others): \[\tan x=\frac{ \sin x }{ \cos x }\]

OpenStudy (anonymous):

thank you :) can you help with this? The equation sin(pi/3-x) is equal to _____?

OpenStudy (mathmale):

If you want to evaluate\[\tan(x-\pi/3)\] you could either memorize and apply the rule for tan (a+b), or you could find sin (x-

OpenStudy (mathmale):

.... Are you finished with this problem? Did you want to move on to sin(pi/3-x)?

OpenStudy (anonymous):

yes i finished the first one, move on please

OpenStudy (mathmale):

Unfortunately, sin(pi/3-x) is not in itself an equation; if there is an equation, where's the rest of it?

OpenStudy (mathmale):

Were you asking to evaluate the expression (not equation) sin(pi/3-x)?

OpenStudy (anonymous):

the question asked what sin(pi/3-x) is equivalent to

OpenStudy (mathmale):

this expression has the form sin (a-b), and you and I have gone over that particular expression before: sin (a-b) = sin a cos b - cos a sin b. Look familiar? if so, what's your next step?

OpenStudy (anonymous):

not quite sure what to do with that

OpenStudy (mathmale):

Hints: \[\sin \frac{ \pi }{ 3 }=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

okay, and then from there?

OpenStudy (mathmale):

cos (pi/3) = 1/2

OpenStudy (anonymous):

alright

OpenStudy (mathmale):

Look at the values for those two trig functions; we obtained them a minute or so back. Substitute those values into the most recent equation.

OpenStudy (mathmale):

You wanted to evaluate 

OpenStudy (mathmale):

and the proper formula to use to do that is\[\sin (a-b) = \sin a \cos b - \cos a \sin b.\]